OpenStudy (anonymous):

a company rented a copy machine for $225 per month for 3 months.the company also paid $0.012 for each copy made.if the total cost of using the copy machine during the 3 month period was $1,107, how many copies did the company make?

OpenStudy (anonymous):

36,000 copies

OpenStudy (texaschic101):

3(225) + .012x = 1107 (x is the number of copies) 675 + .012x = 1107 (subtract 675 from each side) .012x = 1107 - 675 (combine like terms) .012x = 432 (divide both sides by .012) x = 432/.012 x = 36,000 copies JoRily is correct...it is 36,000 copies :)